# ANVIL: Anomaly-based Vulnerability Identification without Labelled Training Data

### v1

@@ III-E1 Line-level Vulnerability Classification
We first evaluate ANVIL against LineVul and LineVD regarding their capability to classify line-level vulnerabilities using ROC-AUC scores. As ANVIL operates differently from the two other tools, we have made the following adjustments.
First, although individual lines serve as the data samples for analysis, both baselines take functions instead of entire programs as inputs. As a result, we utilized the Tree-sitter [29] parsing tool to extract the 158 vulnerable functions that include the 251 vulnerable LOC, from the Magma dataset. Additionally, to mitigate potential biases from only evaluating vulnerable functions, such as characteristic differences between non-vulnerable lines in vulnerable versus non-vulnerable functions, we also selected an equal number of non-vulnerable functions (158) to match the vulnerable ones. This results in a dataset comprising 251 vulnerable LOC and 33,157 non-vulnerable LOC (totaling 33,408 LOC). This approach, while creating an imbalance, is a necessary compromise to simulate a realistic scenario in which we can understand how a detector might perform in practice.
However, LineVul depends on CodeBert for generating embeddings, which can only process up to 512 input tokens. When given a function that exceeds this limit, LineVul truncates the input to the first 512 tokens, thus only evaluating 7,075 out of the total 33,408 lines in the dataset. In contrast, ANVIL evaluates all the 33,408 lines. Additionally, given that LineVul only supports line-level detection within functions it classifies as vulnerable, for functions deemed non-vulnerable, we automatically consider all its lines as non-vulnerable.
Furthermore, LineVD has a different method of determining which lines are considered vulnerable when comparing a CVE with its patched version. Specifically, when a patch to a vulnerability adds new lines, LineVD considers lines that are control- or data-dependent on patched lines as vulnerable lines, with the exception of the added line itself. In contrast, ANVIL would mark the line where the added line should have appeared as the location of the vulnerability. To resolve this discrepancy, we only compare vulnerabilities where the vulnerability fixing patch deletes or modifies lines, as both LineVD and ANVIL treat such lines the same way. After this adjustment, 135 LOC from 76 functions remain classified as vulnerable for this evaluation.
As shown in Table IV, after evaluating the classification performance on ANVIL and LineVul, we observed that ANVIL achieves a ROC-AUC score of 65.25%, which is 1.29 times higher than LineVul’s score of 50.66%. Similarly, compared with LineVD, ANVIL achieved a ROC-AUC score of 72.06%, while LineVD recorded a score of 70.69%. Note that ANVIL’s accuracy is different for LineVul and LineVD due to the different accommodations we had to make to the dataset for each tool. The improvement of ANVIL over LineVD is modest at 102%, yet this result is noteworthy. LineVD’s consistency mechanism enhances its line-level classification by leveraging insights from its function-level classifier, which can reduce false positives in non-vulnerable functions. Despite ANVIL not benefiting from similar function-level support, it still manages to achieve slightly better classification efficacy than LineVD.
TABLE IV: ANVIL vs. Baselines
Baseline ANVIL
LineVul ROC-AUC 50.66% 65.25% (129$\uparrow$)
Top-5 26% 42.05% (162$\uparrow$)
LineVD ROC-AUC 70.69% 72.06% (102$\uparrow$)
Top-5 22.37% 48.68% (218$\uparrow$)
III-E2 Vulnerability Prioritization.
@@
This experiment assesses the ability of the vulnerability detectors to prioritize vulnerable lines within a given vulnerable function by comparing their Top-5 accuracies. Specifically, the Top-5 accuracy metric evaluates whether at least one vulnerable line is present among the top 5 lines ranked by these detectors. For ANVIL, lines are ranked based on the anomaly score, while LineVul uses attention scores, and LineVD employs logits from a MLP for their line ranking. In this experiment, we used the 158 vulnerable functions from Magma as our sample set.
We first compare ANVIL with LineVul. Due to the token limit issue discussed earlier, only 88 of the 158 vulnerable functions have vulnerable lines within the first 512 tokens. Thus, to ensure a fair comparison, both ANVIL and LineVul are evaluated on these 88 functions. The results, presented in Table IV, show that ANVIL significantly outperforms LineVul, with a Top-5 accuracy of 42.05% compared to LineVul’s 26%.
Comparing ANVIL with LineVD on the 76 vulnerable functions, as shown in Table IV, ANVIL achieved a Top-5 accuracy of 48.68%, markedly outperforming LineVD, which managed only 22.37%. This corresponds to an improvement of 2.18 times. Unlike the previous classification task, this evaluation on prioritization focused solely on vulnerable functions, negating the advantage of LineVD’s function-level classifier. Consequently, both tools were assessed under equivalent conditions, leading to the significantly better performance observed with ANVIL in this context.
@@ III-F RQ5: Data Leakage
In this section, we address concerns about potential data leakage that may arise if our patches for the vulnerabilities in our evaluation dataset were present in the LLM training data, as this may enable the LLM to correctly generate the patched code instead of the vulnerable code, as discussed in Section I and III-A. Due to time constraints and the extensive scope of validations required across multiple models, this evaluation is primarily conducted on the CodeLlama-13B model. For this experiment, we continue to use the hybrid scoring function and MCS configuration.
We begin by assessing ANVIL’s ability to detect vulnerable lines as anomalies within the 2024 CVEFixes dataset, which includes 156 vulnerable lines. To ensure a balanced evaluation, we also randomly selected an equal number of 156 non-vulnerable lines from the same source files. The result indicates a P-value of 5.4e-6 between vulnerable and non-vulnerable lines, which is significantly lower than the $\alpha=0.05$ threshold and is comparable to the P-value obtained in the Magma dataset (1.3e-6, from Figure 5). These findings confirm that ANVIL’s vulnerability-anomaly detection capabilities are not tied to dataset exposure and can generalize to unseen vulnerabilities.
Furthermore, to assess the generalizability of ANVIL in detecting line-level vulnerabilities, we evaluated its performance on both our leakage-free 2024 CVEFixes dataset and the entire Magma dataset. Note that earlier experiments with Magma only utilize a subset to accommodate comparison with previous detectors. Similar to previous experiments (section III-E), we evaluate vulnerability classification through ROC-AUC scores. For the CVEFixes dataset, we analyzed a total of 208 functions—comprising an equal split of 104 vulnerable and 104 non-vulnerable functions—from the 2024 CVEFixes dataset. This dataset encompasses 12,402 LOC. The results, shown in Table V, demonstrate that ANVIL maintains consistent ROC-AUC scores across both datasets, achieving 67.20% on the CVEFixes dataset, compared with 65.25% on the Magma dataset. This consistency demonstrates the robustness of ANVIL’s vulnerability detection capabilities across diverse datasets.
For the evaluation of vulnerability prioritization, we focused on the 104 vulnerable functions containing 144 vulnerable lines from the 2024 CVEFixes dataset. To provide a comparative analysis with the Magma dataset, we re-executed the prioritization experiment using all 158 vulnerable functions it contains. Results in Table V reveal that the Top-5 accuracy on the CVEFixes dataset is slightly lower than that observed with Magma. This discrepancy can be attributed to the manual inspection process used for the CVEFixes dataset, detailed in Section III-A, where our limited familiarity with the code bases might have led to some vulnerabilities being overlooked and incorrectly labelled as non-vulnerable. These overlooked vulnerabilities might appear among the top 5 predictions from ANVIL, contributing to the nominally lower Top-5 accuracy. To mitigate the effects of this potential mislabeling, we extended our analysis to include Top-10 accuracy, providing a more lenient criterion that better captures ANVIL’s performance. In this broader evaluation, ANVIL achieved a Top-10 accuracy of 51.92% on the CVEFixes dataset, slightly outperforming the 48.73% observed with Magma, confirming that our anomaly-based approach is robust and not compromised by data leakage in its ability to prioritize vulnerabilities accurately.
TABLE V: ANVIL’s generalizability of Vulnerability Detection
ROC-AUC Top-5 Top-10
Magma 65.25% 34.18% 48.73%
2024 CVEFixes 67.20% 23.08% 51.92%
@@ IV Discussion
IV-A Mechanisms in Line-level Vulnerability Detection
In Section III-E, we evaluated ANVIL for line-level vulnerability detection using the anomaly score (AS) as the foundation, supplemented by two additional filtering mechanisms, M1 and M2, designed to exclude lines unlikely to contain vulnerabilities. To assess the individual contributions of these mechanisms, we conducted an ablation study.
Table VI presents the impact of each component. For clarity and context, we include the baseline performance of ANVIL without any mechanisms and with an anomaly score of 0 for all lines, akin to random guessing, at the bottom of the table. The inclusion of anomaly score alone yielded improvements of 1.12 times in ROC-AUC and 1.1 times in Top-5 accuracy over random guessing. Adding M1 further enhanced the ROC-AUC to 57%, though it did not impact the Top-5 accuracy. However, for Top-10 accuracy, M1’s inclusion increased the accuracy marginally from 43% to 43.7%. When employing all three mechanisms together, there was a substantial increase in both ROC-AUC and Top-5 accuracy, achieving 31% and 32% improvements respectively compared to the baseline.
TABLE VI: Ablation Study
AS M1 M2 ROC-AUC Top-5 Top-10
65.3% (+31$\uparrow$) 34.2% (+32$\uparrow$) 48.7% (+48$\uparrow$)
57.0% (+14$\uparrow$) 28.5% (+10$\uparrow$) 43.7% (+33$\uparrow$)
56.1% (+12$\uparrow$) 28.5% (+10$\uparrow$) 43.0% (+31$\uparrow$)
Random 50% 26.0% 32.9%
IV-B Vulnerable vs Patched lines
In addition to collecting vulnerable and non-vulnerable lines, we also gathered 353 patched lines from the Magma dataset for the 138 known CVEs. These lines include the additions and modifications made to address the CVEs. Additionally, when a defect-fixing commit deletes a line, we also include the subsequent line in the same location as part of the patched lines.
@@ IV-B Vulnerable vs Patched lines
Further analysis using the median anomaly scores under fixed context sizes revealed that patched lines generally align closely with non-vulnerable lines, albeit slightly higher. This trend shifts markedly under the maximum compound statement (MCS) context, where the exact match accuracy for patched lines drops below 50%, causing their median anomaly scores to approach those of vulnerable lines (see Figure 6 right).
Figure 6: Patch vs. CVE vs. Non-vulnerable Lines
To probe deeper, we assessed the statistical significance between patched and other line types using P-values (Table VII). All P-values comparing patched to vulnerable lines fell below the $\alpha=0.05$ threshold, suggesting significant differences. Conversely, the P-values comparing patched to non-vulnerable lines were above this threshold but still displayed relatively low confidence levels (four of six numbers range from 8.36% to 26.67%).
TABLE VII: Patch vs. CVE vs. Non-vulnerable Lines in P-values
Context Size
P-value 100 200 300 400 500 mcs
Patch vs. Vul 0.0040 0.0048 0.0037 0.0035 0.0209 0.0047
Patch vs. Non-vul 0.7944 0.2667 0.1266 0.5102 0.2647 0.0836
The conclusion we draw from these experiments is that even after a vulnerability has been patched, the code that fixes the vulnerability still exhibits some statistical differences with regular code. One potential explanation is that some percentage of patched code is still vulnerable which is consistent with insights from the Magma study [19], so the patched lines are actually a mix of vulnerable and non-vulnerable code. Another potential explanation is that the code locations where vulnerabilities tend to exist may simply have higher entropy, or depend on global factors, which makes it difficult for current LLMs (and possibly their human counterparts) to generate correctly.
@@ V Threats to Validity
Internal Validity. One threat to internal validity lies in data labelling. Correctly labelling which lines are vulnerable continues to be a challenging problem, as vulnerabilities are often difficult to isolate and determining what code is considered vulnerable can take a wide array of definitions. To mitigate this, we use the Magma dataset for our vulnerability detection experiments (RQ1-4). As mentioned in section III, all labelled vulnerabilities in Magma are provided with a POV through extensive manual review and fuzzing campaigns, which ensures that lines labelled as vulnerable are indeed related to the vulnerabilities. Further, manual labelling by a small set of developers, results in better consistency that is less impacted by different definitions of ”vulnerable” that may arise in datasets scraped from CVE reports. For the leakage experiment (RQ5), we manually curate the 2024 CVE-Fixes datasets to ensure the correctness of the labelling. However, while manual curation on its own is likely to have fewer mislabels than simply scraping data from CVE reports, it can still be imperfect due to the complex structure and the large code base of the included projects.
External Validity. While we demonstrated in Section III-F that ANVIL generalizes to non-leakage data by evaluating its consistency across the Magma and 2024 CVEFixes datasets, we concede that numerically comparing the accuracy across two data sets may yield an imprecise conclusion since the two datasets are composed of two different samples of vulnerabilities, and it is nearly impossible to control for this. While we could increase the number of samples in both datasets to decrease chances of sample error, manually finding and reproducing real vulnerabilities is manually intensive and difficult. Future research could aim to collect a more extensive non-leakage dataset to conduct a larger-scale evaluation, providing a more robust statistical comparison of ANVIL’s generalizability to non-leakage data.

### v2

@@ III-E1 Line-level Vulnerability Classification
We first evaluate ANVIL against LineVul, LineVD, and LLMAO in their ability to classify line-level vulnerabilities using ROC-AUC scores. This experiment uses the PrimeVul dataset with several accommodations to ensure fairness and compatibility. First, following the independent testing principle outlined in [nong_vgx_2024, nong_vulgen_2023, chen_diversevul_2023], we ensured that the testing and training sets include data from different sources to avoid bias caused by duplicated data distributions. Specifically, we included only projects not part of the training datasets (BigVul and Devign) used for the baselines. For the remaining data, we included only functions with corresponding source files available to support varying context window sizes used in ANVIL and LLMAO. Furthermore, in line with LineVD and LLMAO, all whitespace lines were removed from the source files. To identify vulnerable lines, we considered only vulnerable functions with paired patched versions, comparing the differences between the two to obtain a superset of potentially vulnerable modified lines. Comments were then filtered out to create the final set of vulnerable lines. This process resulted in a dataset containing 924 vulnerable LoC and 33,820 non-vulnerable LoC from 273 vulnerable functions, along with 325,353 LoC from 9,273 benign functions. These vulnerabilities span 40 different CWEs, with detailed breakdowns provided in Appendix B.
Additionally, as ANVIL operates differently from the three other tools, we made the following adjustments. LineVul depends on CodeBERT for generating embeddings, which is limited to processing a maximum of 512 input tokens. For functions longer than 512 tokens, LineVul truncates them to include only the first 512 tokens, discarding the rest of the function. Consequently, we evaluate LineVul only on lines of code that it keeps, which is 455 vulnerable LoC and 192,382 non-vulnerable LoC in the dataset, whereas the other three tools are evaluated on the full dataset of approximately 360k LoC. Notably, LineVD, another CodeBERT-based detector, does not have this limitation because it generates embeddings for each line individually, allowing it to process longer functions without truncation.
Another limitation of LineVul is that it supports line-level detection only within functions it classifies as vulnerable, whereas line-level detection is not performed for functions deemed non-vulnerable, as per its original implementation. In contrast, all other tools support line-level detection across all functions. To maintain consistency in evaluation, we adopted the method used in LineVD: for functions classified as non-vulnerable by LineVul, all lines within those functions are automatically considered non-vulnerable.
Furthermore, LineVD employs a different method for identifying vulnerable lines when comparing a CVE with its patched version. Specifically, in addition to lines directly modified in the patch, LineVD considers lines that are control- or data-dependent on the patched lines as vulnerable. In contrast, ANVIL and the other two baselines label only the patched lines as vulnerable. To resolve this discrepancy and ensure a fair comparison, we disabled LineVD’s dependency mechanism during our experiments.
As shown in Table IV, ANVIL achieves ROC-AUC scores of 56.8% and 55.7% with the CodeLlama-7B and 13B models, respectively, outperforming all three baselines. Specifically, ANVIL (13B) achieves 1.10 times higher ROC-AUC than the best configuration of LLMAO (16B) and 1.15 times higher than both LineVD and LineVul. Notably, LineVul and LineVD achieve only 49.6% ROC-AUC, which is essentially the same as random guessing (50%), while LLMAO performs marginally better than random. Moreover, when comparing different LLM sizes within the same model (e.g. 7B vs. 13B), the improvement in performance is modest, with gains of less than 1.02 times, as observed in both ANVIL and LLMAO.
III-E2 Vulnerability Prioritization
@@
This experiment evaluates the ability of vulnerability detectors to prioritize vulnerable lines within a given vulnerable function by comparing their Top-N accuracies and Mean First Ranking (MFR). Specifically, the Top-N accuracy metric measures whether at least one vulnerable line is included among the top N lines ranked by each detector. For ANVIL, lines are ranked based on their anomaly scores, while LineVul ranks lines using attention scores, and LineVD and LLMAO rely on logits from their respective neural network heads for line ranking. We report Top-1, Top-3, and Top-5 accuracies, as these metrics are originally used by the baselines. Additionally, we calculate MFR (smaller is better) by averaging the rank of the first vulnerable line in each function, providing a comprehensive measure of the tools’ overall vulnerability prioritization performance. This experiment uses 273 vulnerable functions from the PrimeVul dataset, consisting of 924 vulnerable LoC and 33,820 non-vulnerable LoC.
As shown in Table IV, ANVIL achieves significantly higher Top-N accuracies compared to its baselines. While ANVIL achieves a moderate 1.12 times better Top-1 accuracy than the best-performing version of LLMAO (16B) and LineVD, its performance becomes more pronounced when the criterion is relaxed to Top-3 and Top-5, with ANVIL achieving 1.40 and 1.41 times higher accuracies, respectively. Notably, the 512-token limit of LineVul results in an average of only 25 available LoC per function, compared to an average function length of approximately 141 LoC for the other detectors. This shortened context theoretically gives LineVul an advantage in achieving higher Top-N accuracy due to the smaller candidate pool. However, its performance falls short, achieving only 4.0% Top-1 accuracy (the lowest among all detectors), and its Top-3 and Top-5 accuracies remain significantly below those of ANVIL.
TABLE V: ANVIL’s generalizability on 2024 CVEFixes Dataset
[…]
ANVIL (13B) PrimeVul 9.9% 27.8% 37.4% 53.1% 26.8 0.22 56.8%
ANVIL (13B) 2024 CVEFixes 9.5% 26.5% 40.1% 67.4% 16.6 0.22 67.9%
This shortened function length also explains why LineVul is biased in achieving the highest MFR. To provide a more objective and unbiased comparison, we introduce the Normalized Mean First Ranking (N-MFR), which adjusts the MFR by normalizing it against the average function length. For LineVul, we apply a normalization factor of 25 LoC, while for all other detectors, we use an average function length of 141 LoC. The N-MFR results reveal that ANVIL with CodeLlama-13B achieves the best score of 0.22, which is 44% lower than LLMAO (16B) and 27% lower than LineVD. Although LineVul originally records the best MFR, its normalized version drops to 0.71, making it the worst performer among all detectors.
The impact of LLM size on detection performance appears to be limited. For ANVIL, using the larger CodeLlama-13B model does not result in significant improvements over the 7B version. Interestingly, for Top-1 accuracy, the 7B model slightly outperforms the 13B model by 0.4%. This trend is even more pronounced with LLMAO, where the 6B version generally outperforms the 16B version, except at Top-1 accuracy. These observations align with our earlier findings on ROC-AUC, suggesting that smaller models can sometimes offer better performance and may be a more practical choice than their larger counterparts.
@@ III-F RQ5: Data Leakage
In this section, we address concerns about potential data leakage that may arise if our patches for the vulnerabilities in our evaluation dataset were present in the LLM training data, as this may enable the LLM to correctly generate the patched code instead of the vulnerable code, as discussed in III-A. Due to time constraints and the extensive scope of validations required across multiple models, this evaluation is primarily conducted on the CodeLlama-13B model. For this experiment, we continue to use the hybrid scoring function and AC configuration.
We assess the generalizability of ANVIL in detecting line-level vulnerabilities by comparing its detection performance between the aforementioned PrimeVul dataset and our newly collected 2024 CVEFixes dataset, as described in III-A. Similar to RQ4, this evaluation examines both vulnerability classification and prioritization performance.
Similar to previous experiments, we evaluate vulnerability prioritization using Top-N accuracies, MFR, and N-MFR. For N-MFR, we apply a normalization factor of 76 LoC, representing the average vulnerable function length in the 2024 CVEFixes dataset. The evaluation is conducted on the 147 vulnerable functions extracted from 100 CVEs. As shown in Table V, ANVIL demonstrates consistent Top-1 to Top-5 accuracies across both datasets. However, while these accuracies remain at similar levels, we initially expected slightly higher metrics for CVEFixes compared to PrimeVul due to the shorter average function length in CVEFixes. This discrepancy possibly arises from the manual inspection process described in Section III-A, where our limited familiarity with the codebases may have resulted in some vulnerabilities being overlooked and mislabelled as non-vulnerable. These overlooked vulnerabilities could appear among ANVIL’s top predictions, leading to nominally lower Top-1 and Top-3 accuracies.
To address the potential effects of this mislabelling, we extended our analysis to include Top-10 accuracy, providing a more lenient criterion to better reflect ANVIL’s performance. Under this expanded evaluation, ANVIL achieved Top-5 and Top-10 accuracies of 40.1% and 67.4%, respectively, on the 2024 CVEFixes dataset, outperforming the corresponding 37.4% and 53.1% accuracies observed with PrimeVul. This aligns with our earlier claim that ANVIL should theoretically perform better on the CVEFixes dataset due to its shorter function lengths. Notably, as the criterion is relaxed from Top-5 to Top-10, the performance gap between the two datasets widens further, highlighting ANVIL’s ability to prioritize vulnerabilities effectively within the top-ranked lines in a function. Additionally, after normalization, the N-MFR for both datasets converges to a score of 0.22. This consistency underscores the robustness of ANVIL’s vulnerability prioritization capabilities, even when applied to unseen data.
For the evaluation of vulnerability classification, we assess all 3,736 functions in the 2024 CVEFixes dataset, which consists of 235 vulnerable LoC and over 110k non-vulnerable LoC, using ROC-AUC scores. As shown in Table V, ANVIL achieves a surprisingly higher ROC-AUC score of 67.9% on the 2024 CVEFixes dataset compared to 56.8% on the PrimeVul dataset. This result confirms that our anomaly-based approach is robust and remains unaffected by potential data leakage, accurately detecting vulnerabilities even in unseen datasets. The observed difference in ROC-AUC scores may be influenced by various factors, such as differences in average function length, dataset difficulty, the degree of sparsity of vulnerabilities, or the ratios between vulnerable and benign functions. Future research could systematically explore how these dataset characteristics impact the performance of vulnerability detectors.
@@ IV Discussion
IV-A Vulnerable vs Patched lines
TABLE VI: Patch vs. Vul vs. Non-vul Lines in ROC-AUC
Context Size
P-value 2 10 20 40 60 80 100 200 300 400 500 AC
Patch vs. Vul 53.4% 53.9% 54.0% 55.5% 55.3% 56.1% 55.4% 57.1% 57.1% 57.1% 57.2% 56.8%
Patch vs. Non-vul 50.8% 51.1% 50.7% 52.3% 53.7% 53.9% 54.0% 52.7% 54.8% 53.7% 51.7% 55.2%
In addition to collecting 512 vulnerable and non-vulnerable lines, we also gathered 353 patched lines from the Magma dataset for the 138 known CVEs. These lines include the additions and modifications made to address the CVEs. Additionally, when a defect-fixing commit deletes a line, we also include the subsequent line in the same location as part of the patched lines.
@@ IV-A Vulnerable vs Patched lines
Further analysis using the median anomaly scores under fixed context sizes revealed that patched lines generally align closely with non-vulnerable lines, albeit slightly higher. This trend shifts markedly under the Adaptive Context (AC) settings, where the exact match accuracy for patched lines drops below 50%, causing their median anomaly scores to approach those of vulnerable lines (see Figure 7 right).
Figure 7: Patch vs. Vulnerable vs. Non-vulnerable Lines
To explore this further, we evaluated our approach’s ability to distinguish patched lines from other line types using ROC-AUC scores, as shown in Table VI. All scores comparing patched lines to other line types exceeded the 50% threshold of random guessing, indicating that ANVIL differentiates among the three line types. However, the extent to which the scores surpass the threshold varies. Specifically, the ROC-AUC scores for distinguishing patched lines from vulnerable lines exceed the threshold by more than 1.14 times, whereas the scores for differentiating patched lines from non-vulnerable lines show a narrower margin, ranging from 50.7% to 55.2%.
The conclusion we draw from these experiments is that even after a vulnerability has been patched, the code that fixes the vulnerability still exhibits some statistical differences with regular code. One potential explanation is that some percentage of patched code is still vulnerable which is consistent with insights from the Magma study [hazimeh_magma_2020], so the patched lines are actually a mix of vulnerable and non-vulnerable code. Another potential explanation is that the code locations where vulnerabilities tend to exist may simply have higher entropy, or depend on global factors, which makes it difficult for current LLMs (and possibly even for humans) to generate correctly.
@@ V Threats to Validity
Internal Validity. One threat to internal validity lies in data labelling. Correctly labelling which lines are vulnerable continues to be a challenging problem, as vulnerabilities are often difficult to isolate and determining what code is considered vulnerable can take a wide array of definitions. To mitigate this, we use the Magma dataset for our vulnerability detection experiments (RQ1-3). As mentioned in Section III, all labelled vulnerabilities in Magma are provided with a PoV through extensive manual review and fuzzing campaigns, which ensures that lines labelled as vulnerable are indeed related to the vulnerabilities. Further, manual labelling by a small set of developers, results in better consistency that is less impacted by different definitions of “vulnerable” that may arise in datasets scraped from CVE reports.
Maintaining a diverse range of code samples while ensuring high labeling accuracy is always a trade-off in generating vulnerability datasets. To the best of our knowledge, we choose the PrimeVul dataset in RQ4 due to its extensive diversity of open-source projects and the large number of vulnerabilities it contains. Moreover, as mentioned earlier, PrimeVul validates the authenticity of vulnerabilities by automatically analyzing corresponding vulnerability descriptions from the NVD database. This process enables PrimeVul to achieve a labeling accuracy of over 90%, outperforming other vulnerability datasets such as Devign [zhou_devign_2019], DiverseVul [chen_diversevul_2023], and CrossVul [nikitopoulos_crossvul_2021].
For the leakage experiment (RQ5), we manually curate the 2024 CVE-Fixes datasets to ensure the correctness of the labelling. However, while manual curation on its own is likely to have fewer mislabels than simply scraping data from CVE reports, it can still be imperfect due to the complex structure and the large code base of the included projects.
External Validity. While we demonstrated in Section III-F that ANVIL generalizes to non-leakage data by evaluating its consistency across the PrimeVul and 2024 CVEFixes datasets, we concede that numerically comparing the accuracy across two data sets may yield an imprecise conclusion since the two datasets are composed of two different samples of vulnerabilities, and it is nearly impossible to control this. While we could increase the number of samples in both datasets to decrease the chances of sample error, manually finding and reproducing real vulnerabilities is intensive and difficult. Future research could aim to collect a more extensive non-leakage dataset to conduct a larger-scale evaluation, providing a more robust statistical comparison of ANVIL’s generalizability to non-leakage data.